This Wiki it is a logbook about high power targetry

An experience called IS 1149 at TRIUMF results from a collaboration between CERN and TRIUMF. The objective of this experiment is the production of high intense 17-25 Ne beams with a 0.5 GeV proton beam with a sigma of 2.6 mm and a maximum beam intensity of 50muA. The target used in this experience is an alumina - Nb composite material developed at ISOLDE-CERN within the EURISOL task#3.

The schedule for all TRIUMF experimental areas: http://mis.triumf.ca:10040/schedule/view

The ISAC operations logbook: https://isac.triumf.ca/ISAC_Operations_Elog/frames.pl

The revised and updated schedule (5th February) : http://www.eurisol.org/site02/meetings.php?nID=292&page=1

Planning the dates and missions: (as proposed by Thierry Stora 17.03.2009)

1) Stable beam: look for AlF, AlO, Al, O, O2, CO, CO2, C, N2, alkaline, earth-alkaline ---> 18 March The target is very clean! No significant Al and O have seen, K and Ca detected.

2) Source efficiency measured with a Ne and Ar gas in function of the cathode temperature and magnet ---> 19 - 20 March? Comparison with the measured efficiencies at ISOLDE with a hot FEBIAD. 18/03/2009 Problem with the synchrotron RF cavity frequency 19/03/2009 morning Ne flow issues in ITE have been resolved. Now the target heater is being raised to 580A (ion source at 280A). Mass scan running overnight. He and Kr gas leak are not available; we can only count with Ne and Xe ones (Pierre) Pierre got an efficiency for Ne of 1E-5, he thinks that the efficiency value is wrong. 20/03/2009 Mass scan revealed the existence of O, K.

3) Starting of the proton beam irradiation with 1muA current and a target Imax 580A (approximately 1100C) ---> 21 March i) Measurement of the yields for all Ne isotopes (17 to 25) ii) Measurement with a Faraday cup the radioisotopes stable Ne beams (20, 21 and 22). In the range of 10pAs?? iii) Measurement of He, F, Kr, molecules (tricky: mixes of CO and N2 and all unimaginable combination; CO2 will be better). Note: gamma yield and T1/2 measurements needed. 21March - 1 uA proton beam hitting the target since 8:00 AM; total resistance in the target is 1.2E-2 ohm and didn't change significantly after proton beam, which according to Etam would correspond to a temperature increase of 40C with a good contact or 80C with a bad contact. According to the table listed at Fig. 1, the temperature increase in the resistance was 13 deg C. The total beam current in the target was around 1.5E-10A before turning on the proton beam and after 3.0 E-9 A, the difference is due to the radioisotope species. Protons gone on and off, water cooling of the pumps in the high radiation zone is giving an alarm, target cooled down, no current on the target. Until this problem is solved no other actions can be taken. No source efficiency data taken yet. Problem solved at 8pm. At around 11pm, 19F seen in the mass scan, a light in the end of the tunnel, after 3 days of impatient waiting.

4) Proton 3 muA. ---> 22 March i) Verification of impurities and their effect in the yields (for example Ne series, Ne source efficiency, modulation of the external current) We get 1 muA proton since early morning 7am or so.

Pierre had tuned ion source (Cathode and magnet). p= 2.5e-6mbar. We have seen 18Ne and 19Ne. ca 5e2 18Ne/s and 2.2e3/s 19Ne . Sigma proton =9-10 mm. (ABRABLA in-target prod 16g/cm2: 5.3e8/muC for 18Ne. With eion=1% and etarget=1% this gives 5e4/muC: we miss x100 or so, but target rather cold)

Now they will tune the beam on the target. reduce FWHM to 4.5-5 mm Can get up to 8e3/s 19Ne. Have found beam slightly tilted and have corrected it. Further tuning (x25 missing)

Marik had on SiC HP target with hot plasma (15g/cm2): 70muA proton : 7e7pps 19Ne ==> 1e6/muC 8e6pps 18Ne ==> 1e5/muC 2e5pps 17Ne ==> 3e3/muC

Sandrina on shift until 9pm, Etam dropped at the airport at 1pm, Thierry takes over at 9pm until Monday morning.


P1000829.JPG

Plan: further tuning with 1muA. then full Ne yields. Then raise up to 5muA with 1 or 2 steps: Ne series, and all other kind of yields.

The estimated yields are: 18Ne 2.8E2/s 19Ne 9.3E3/s 23Ne 7.9E3/s 24Ne 2.0E3/s

Proton beam is now 3 muA since 17:15 . The resistance of the target increased from 0.0118381 to 0.011918 ohm. The mass 16 was checked to see if there's any variations on the oxygen levels. After a few minutes it still increasing, let's see later if this effect it is permanent (from 230pA to 400pA). A picture is indexed with proton beam current, resistivity, current for 16O, power.

The estimated yields are now: 17Ne 11/s 18Ne 1.3E3/s 19Ne 2.4E4/s 23Ne 2.2E3/s 24Ne 5.6E3/s

Thierry: Protons at 5muA since about 10pm. Yield estimates: 17Ne 25/s 18Ne 3.7E3/s 19Ne 1.0E5/s 23Ne 7.0E4/s 24Ne 1.4E4/s

Target resistivity has gone from XXX to 0.01196 ohm. This should correspond to container at ca

At ca midnight, Pierre is tuning the ion source on 18Ne. yield improved x2, ca 160nA total current, small drain current 30mA. Cathode quite cold, magnet increased up to 100A.

5) Proton 5 muA. ---> 23 March i) Verification of impurities and their effect in the yields (for example Ne series, Ne source efficiency, modulation of the external current)

Protons were lost for a short while. this has happened once before at 5muA. Yields after ion source optimization: 18Ne : 5.2e3pps

1:15 am Decrease target heating to 520A.

18Ne: 4.7e3pps

1:30 am proton intensity at 7.5pmuA 2:30 am measure Ne: FWHM = 6.1mm

17Ne 92/s

18Ne 1.1e4/s (done at 2am)

19Ne 3.5e5/s

A=20 stable: 2.8e8pps

A=21 2.1e7pps

A=22 1.4e8pps

23Ne 2.1e5/s (2.5nA stable)

24Ne 5.4e3/s; repeated and confirmed

4:30am container resistance at 0.011800 ohm

5:15am

repeated 18Ne 1.1e4/s : good , we are running under stable conditions.

Doing mass scan A=1-210 amu.

7am: start cool down target at 400A (stable protons at 7.5muA overnight), to increase proton beam to 10muA.

8 am : target resistance 0.010945 ohm. Yield of 18Ne:5.8e3pps (7.5muA) protons at 10muA for a short while, then protons stopped. Protons tripped a 2nd time.

Measuring the Ne isotope series:

17Ne 97/s

18Ne 1.4e4/s

19Ne 4.4E5/s

A=20 stable: pps

A=21 pps

A=22 pps

23Ne 3.4E5 /s (nA stable)

24Ne 1.3E4/s;

Ri=0.0112322 ohm 11:20 Ohmic heating changed, we have now in the ion source 311.6, 3.009V and in the target 360A, 3.94V. Rf=0.010321 ohm

17Ne 84/s 18Ne 1.5E4/s

12:00 R=0.0110209 ohm

12:00 no protons due to intervention in cyclotron. Work out gas leak and find about ion source efficiency during the afternoon; Afterward, resume with ramping up proton current to reach 20muA.

17:58 R=0.0110412 ohm

18:10 Now at 12.5 muA p+, FWHM 6.6 mm.

18:48 R=0.0111472 ohm

17Ne 1.7E2/s 18Ne 2.1E4/s 19Ne 5.7E5/s 23Ne 6.3E5/s 24Ne 8.4E4/s

Protons off since 20:24 for about 6 min.

25Ne 17/s

There was no protons for about 30min since 21:30. There's a high current for mass 28 stable of 25nA and we don't see 15C16O+. Starting to look into Carbon isotopes yields, if no other proton sparking arrives. Checked mass 15C but it was not seen in the B+ scintillator. It must be other molecule/specie.

Target current down to 300A.

22:22 Raising the proton beam intensity until 15muA.

R=0.01110055 ohm A=16, O current approximately 960pA.

target: 300.7A, 3.14V, source: 311.8A, 3.004V

Estimated yields: 17Ne: 2.4E2/s 18Ne: 3.1E4/s 19Ne: 3.7E5/s

A=18 420 pA A=19 120 pA A=20 stable: 76pA A=21 pps: 10 pA A=22 pps: 25 pA

24/03/2009

Thierry/Aurelia/Peter (00:00 - 8:00 AM)

midnight 19Ne: 7.9e5/s (suspicion that previous value was not with right scintillator) 23Ne: 2.2e6/s 24Ne: 1.2e5/s

Have checked 25Ne, and it is not clear if the count rate seen by Marik is indeed Ne or not.

But we have seen Na:

21Na : 3.0e4/s 25Na : 3.9e4/s 26Na : 4.8e1/s

Take target heating down to 160A (r=????)

18Ne: 3.0e4/s

Raise proton intensity up to 17.5 muA. But a lot of tripping ( at least 5 times) : I hope the target survives.

Actually, cyclotron could not ramp up to 17.5 muA. Went back to 15muA. Check 18Ne: 2.9e4/s

Good.

Keep proton current at 5muA otherwise continuously stripping.

Spent a least an hour to scale the magnet to 74Kr (took us 1 hour to figure out stable Xe++ , but we have an excuse: It's 7am)

From 8:00 to 17:00 intervention in the cyclotron.

Sandrina / Pierre Bricault (5:00-12 PM)

322.8A, 3.131 V FEBIAD 400A Target FWHM H6.3, V5.6

Pierre tuned the FEBIAD and determined the ion source efficiency with a Ne leak, found it to be 0.04. This is low and the drain current does not change from a low value of 31mA.

(On offline pumpstand efficiency was 1.1%, ISOLDE MK7 online is 0.2%-0.8% in database, VD7 is 7% - comment added by Thierry)

Protons at 2 muA

18Ne - 410/s

Protons at 5muA

18Ne - 7.0E3/s

Protons at 10muA

18Ne - 4.1E4/s

Protons at 15muA

Target 300.7A, 3.31v / FEBIAD 322.8A, 3.131v

18Ne - 7.0E4/s

Target cooling... The ohmic current is now 180A.

21:00 Protons at 17.5muA

R=1.064E-2 ohm

17Ne 7.1E2/s 18Ne 7.0E4/s 19Ne 1.7E6/s 23Ne 2.6E6/s 24Ne 5.1E4/s

March 26th

Marik/ Friedhelm (00:00 - 8:00)

Went to 20muA protons, measured Ne series, He, Ar, Kr:

A=27: 1.5nA

17Ne: 4.9E3/s 18Ne: 1.0E5/s 19Ne: 1.9E6/s 20Ne: 1.8E8/s 23Ne: 1.8E6/s 24Ne: 6.4E4/s 25Ne: 1.6E2/s

75Kr: 7.4E4/s 74Kr: 5.4E3/s 73Kr: 1.5E2/s

6He: 8E5/s 8He: 2.7E3/s

35Ar: 1.2E4/s 34Ar: 5e2-7E3/s (betas/gammas)

cyclotron stop during the day

Pierre had a look on the FEBIAD, the drain current is still at 32mA or so. Cathod current 325.9 A (3.43V), DeltaV=197V, coil 100A. 0.2% efficiency on Ne1+

Thierry/Peter on shift

5pm: resume with 5muA protons. Target heating at 400A 4.37V

Check 18Ne :

400A, 5muA: 7E3/s 400A, 10muA: 6.2E4/s 100A, 10muA: 4.6E4/st

A lot of trips to reach 15A, and sometimes no ramping to get back from 0 to 11-13A

100A, 15muA : 1.1E5/s (FWHM=7mm)

100A, 20muA :1.5E5/s

27/03/2009

Sandrina/Pierre (00:00-8:00)

22.5 muA since 01:42!

25Ne - 2E4/s 24Ne - 2E4/s 23Ne - 4E6/s 19Ne - 1.4E6/s 18Ne - 1.8E5/s 17Ne - 1.3E3/s

8He - 3.2E3/s 8Li - 3.5E3/s 6He - 2.6E6/s Good...!

Going now to 25Ne again? Now 25 muA!!!

5:40 25muA

8He - 3.2E3/s 8Li - 3.5E3/s 6He - 1.3E6/s frown

17Ne - 1.2E3/s frown

After retuning the FEBIAD

17Ne 1.61E3/s Good! 18Ne 6.1E4/s

Now retuning the beam...

18Ne 2.6E5/s 19Ne 5.8E6/s 23Ne 8.4E6/s 24Ne 8.2E5/s

9: 45 cyclotron stop during the day

Marik / F. (8:00-12:00PM)

Remeasuring the Ne series for 25muA p+

18Ne - 1.6E5/s 19Ne - 3.0E6/s 23Ne - 6.5E6/s 24Ne -
https://isac.triumf.ca/ISAC_Operations_Elog/frames.pl

INITIAL plans

6) Proton 10 muA. ---> 22 March i) Verification of impurities and their effect in the yields (for example Ne series, Ne source efficiency, modulation of the external current) ii) Beam stability

7) Proton 20 muA. ---> 23 March i) Yields measurements ii) Stability control iii) target container temperature

8) Proton 25 muA. ---> 24 March i) Yields measurements ii) Stability control iii) target container temperature

9) Proton 25 muA. ---> 25 March i) ion source optimization ii) Stability

10) Proton 25 muA. ---> 26 March Physics users if any.

11) Proton 25 muA. ---> 27 March Tests with increase of intensity by 5muA steps (impact on the yields, impurities, target temperature) If not conclusive keep at 25 muA.

12) Proton 25 muA. ---> 28 - 29 March Monitoring of stability.

13) Proton 25 muA. ---> 30 March Wobbling tests to see if it permits to increase the proton beam intensity.

14) End of tests ---> 8:00 AM, 31 March.

Important considerations:

1) If the proton beam is stopped, it must be given a protocol to start slow increase of proton current, for avoiding the full intensity in one goal!

2) The principal isotopes to be measured are Ne isotopes, which would allow us to test the performance of the target and ion-source with the time.

3) This experiment logbook must be updated regularly until the end of the experiment, so if you want to know what is going on with our beautiful Al2O3/Nb target stay connected. Anyone is welcome to had new ideas or comments to this list...and don't forget to add the date from time to time /Richard

26/03/2009 Arrival of Richard and departure of Thierry

Evening spent doing target tests at 25uA. A lack of CO coming through indicates a cool cathode. Target still holding well

27/03/09 Protons off for ion source optimization. Strange phenomena observed while increasing the coil current. There is a definate shift in the beam but after further investigations it turns out to be acceptable within the hall probe tolerance. This took most of the morning.

14:00 Slight but lengthy tuning of the ion source to increase the efficiency by about 50%. Note that we do not know the leak rate of the Ne leak thus the eficiency is not clear yet.

16:00 Decide to do a proton beam scan at 1 uA and definitaly see a shift in best beam position. This is nicely logged in their logbook which I shall try and retrieve. The tune of the beam was determined to be poor but acceptable for our purposes. Beam increased to 25uA for target tests

18:30 Target tests commence. Note that beam size is now ~ 8mm FWHM compared to 6.5 earlier in the week. Tests done on 17 - 25 Ne. Slight gain in yield (~40%) except for 18Ne and 19Ne. This is probably due to the ion source tuning.

21:00 Tests finished. Decided to leave beam at 25uA overnight. Slight increase in vacuum pressure (2.2E-06 mbar) observed but deemed acceptable. I refused the optio of going to 30uA as TITAN would like to take beam tomorrow.

Plans for saturday: Make count rate measurement of 19Ne at 30kV for reference, Decrease and scale separator to 20kV (required by TITAN) and re-do count rate measurement. Keep protons at 25uA and occasionally follow count rate throughout day. Have also asked for integration of protons on target (but not sure that their proton integrator works ...sounds familiar)

28/03/09

09:08 Sandrina arrives at airport for departure at 10:00 (a bit tight)

10:00 Arrive at Triumf to find that the HT has been cut at circa 09:30 due to a water heating of target circuitry. Cause not known but could account for last night's vacuum increasing. Decided to close the water cooling circuitt of the second target station to increase flow rate in our target. The two target station use the same water circuit in parallel. This involves calling in someone to enter the target zone after a certain amount of radioactive cooling.

Its now 14:30 and we're still waiting. As soon as we come back on line we'll resume last night's micro planning. Will monitor count rate today and tomorrow while TITAN run. Will ask for 30uA on Monday and re-do Ne measurements.

14:30 Water circuit modified and protons back. The target is still in good shape confirmed by the count rate measurement of 17Ne being comparable to that of last night (after ion source tuning). Fredheim is currently setting up the separator for 15kV for TITAN.

15:30 Found the proton integrator. At present ~1565uA hours have been taken on target or ~60 hrs at 25uA.

29/03/09 09:00 1950 uA hrs of beam on target. TITAN have given up taking the beam due to impurities in their RFQ. Leave 25uA of beam on target for the next 24 hours. Target tests to resume tomorrow.

30/03/09 09:00 Start tuning beam to tape station. Plan today: push the target and ion source parameters while following yields.

Tests on 30-03-09.

Starting at 10:00

Setting 1 Cathode I 334.8 A U 3.65 V

Target I 100.2 A U 1.08 V

Anode I 0.074 A U 197 V

Magnet I 98 A U 1 V

Puller I 879 uA U 816 V

HT 110 15 kV

Pressure 2.90E-06 mbar 3.20E-06

P-beam 25 uA Observation

17Ne 2.20E+03 cts/s FC14 18Ne current increased from 100pA to 1.2nA

18Ne 3.40E+05 cts/s 23Na measured at 500pA

19Ne 4.80E+06 cts/s p-beam average FWHM 6.6mm

23Ne 5.80E+06 cts/s More beam on mass 23.

24Ne 1.20E+06 cts/s

25Ne 400 - 1200 cts/s

15:30 move target temp to 200A while monitoring 17Ne count rate. Absolutely no change.

Total current 2.7 E-07 A

Cathode I 334.8 A U 3.66 V

Target I 200 A U 2.2 V

Anode I 0.077 A U 197 V

Magnet I 98 A U 1 V

Puller I 879 uA U 816 V

HT 110 15 kV

Pressure 2.70E-06 mbar

1.4E+04 cts/s for 21Na

16:30 start ramping up to 360A. According to the calibrations we have, this would mean a temperature of ~1630C. We would like to monitor the Na21 while heating to have an indication of the cold line temperature. Total beam current at 3.0E-07A

Cathode I 334.9 A U 3.66 V

Target I 260 A U 2.9 V

Anode I 0.079 A U 197 V

Magnet I 98 A U 1 V

Puller I 879 uA U 816 V

HT 116.27 15 kV

Pressure 2.84E-06 mbar

1.4E+04 cts/s for 21Na so no change

17:00 Start ramping target up to 300A….1A/10s

17:05 Protons dropped for 5 minutes. They have difficulty in going beyond 23.5uA

Cathode I 334.9 A U 3.66 V

Target I 300 A U 3.39 V

Anode I 0.081 A U 197 V

Magnet I 98 A U 1 V

Puller I 879 uA U 816 V

HT 117.19 15 kV

Pressure 2.84E-06 mbar

Total current 3.25 E-07A

1.8E+04 cts/s for 21Na

no change for 17Ne.

18:15 Increase target heating to 360A

Cathode I 334.9 A U 3.66 V

Target I 360 A U 4.13 V

Anode I 0.085 A U 197 V

Magnet I 98 A U 1 V

Puller I 879 uA U 816 V

HT 122.99 15 kV

Pressure 3.21E-06 mbar

Total current 3.74 E-07A

2.2E=04 cts/s on 21Na

18:30 Ramping up to 400A

Cathode I 334.9 A U 3.66 V

Target I 400 A U 4.62 V

Anode I 0.088 A U 197 V

Magnet I 98 A U 1 V

Puller I 879 uA U 816 V

HT 122.99 15 kV

Pressure 3.30E-06 mbar

Total current 3.09 E-07A 2783.74uAhrs so far

2.7E+04 cts/s for 21Na

2E+04 cts/s for 17Ne

19:00 Start ramping down the target heating to 100A in preparation to take 27.5uA of p-beam. 19:25 Target at 100A

Cathode I 334.9 A U 3.66 V

Target I 100 A U 1.08 V

Anode I 0.074 A U 197 V

Magnet I 98 A U 1 V

Puller I 879 uA U 816 V

HT 105.29 15 kV

Pressure 3.56E-06 mbar

Total current 2.65 E-07A

19:27 protons dropped

19:28 27.5uA of p-beam on target.

3.3E-07A of total beam

17Ne yield doesn't really increase. 2E+03 cts/s

19:50 P-beam increased to 30uA

17Ne measured at 2.9 E +3cts/s

20:00 Measurement of stable 20Ne for efficiency measurements Ne efficiency estimated at 0.4%. Time to increase the cathode and improve on the efficiency.

22:00 2E+3 cts/s on 17Ne

22:17 Started automatic mass scan. Will take about 2 hours

Here's a resume of what happened overnight

The mass scan ran until ~01:00 after which the cathode was increased to 340A to let it outgas. At the same time the Ne gas inlet valve was opened and left open to monitor Ne. Consequently the pressure in the target increased to 1.2 E -05 mbar. The operators then dropped the p-beam to 20uA for 1 hour until the pressure got back to ~6E-06 mbar. It was left as such until 08:00 this morning. After realizing all this just now, I have asked for the gas valve to be closed in order to get a Ne efficiency, however it will be measured without the contribution of the incidental target heating from the p-beam.

Integration of 3174uAhrs on target...and its still going well! Congratulations to the "EURISOL Team"

The target also withstood 30uA of p-beam (15kW) for about 12 hours. Only the last measurement of 17Ne with a decrease from 2.9E3 to 2 E+03 cts/s gave some cause for concern but this is still within the error bars.


Note on energy deposition: FLUKA vs SRIM

SRIM results: Nb: 1.500 MeV/mm 15.00 MeV/cm Al2O3: 0.628 MeV/mm 6.28 MeV/cm Ti: 0.830 MeV/mm 8.30 MeV/cm

Target thicknesses: Nb = 35 x 0.5 mm = 1.75 cm Al2O3 = 70 x 1.07 mm = 7.5 cm Ti = 70 x 0.1 mm = 0.7 cm

SRIM energy deposited Nb = 26.25 MeV Al2O3 = 47.1 MeV Ti = 5.81 MeV

SRIM Total = 70.16 MeV SRIM: 15.83% total beam (500MeV H-)

FLUKA: 16.22% total beam


Topic attachments
I Attachment History Action Size Date Who Comment
Compressed Zip archivezip All_Chapters_21_June_2010_Dr.Gauckler.zip r1 manage 10446.8 K 2010-06-21 - 17:19 UnknownUser Thesis
PDFpdf All_Chapters_24_June_2010.pdf r1 manage 11484.8 K 2010-06-27 - 18:59 UnknownUser another version
PDFpdf All_Chapters_28_June_2010.pdf r1 manage 11372.9 K 2010-07-03 - 23:45 UnknownUser  
PDFpdf All_Chapters_2_July_2010.pdf r1 manage 11530.3 K 2010-07-03 - 23:24 UnknownUser  
PDFpdf All_Chapters_4_July_2010.pdf r1 manage 11405.3 K 2010-07-07 - 00:39 UnknownUser  
Unknown file formatdocx Chapter_1_Chapter_2.docx r1 manage 2744.9 K 2010-07-16 - 18:14 UnknownUser  
PDFpdf Chapter_3.pdf r1 manage 3677.8 K 2010-07-25 - 21:34 UnknownUser PhD 25 July 2010
PDFpdf First_proton_beam.pdf r1 manage 37.1 K 2009-03-22 - 19:22 SandrinaFernandes  
JPEGjpg P1000829.JPG r1 manage 2803.8 K 2009-03-24 - 01:25 SandrinaFernandes  
PDFpdf R_vs_T.pdf r1 manage 49.7 K 2009-03-22 - 08:25 SandrinaFernandes Target_temperature_control
Unknown file formatpptx S1149_Target_Temperature_Control.pptx r1 manage 102.1 K 2009-03-23 - 07:27 SandrinaFernandes S1149_temperature_control
PDFpdf T_vs_I-protons.pdf r1 manage 30.7 K 2009-03-22 - 19:32 SandrinaFernandes  
PDFpdf Triumf_Target_final.pdf r1 manage 5174.0 K 2009-03-23 - 11:13 ThierryStora Stefano's target dev
Microsoft Excel Spreadsheetxls Yields.xls r5 r4 r3 r2 r1 manage 71.5 K 2009-03-25 - 17:56 ThierryStora Yields during tests, to be updated regularly
PNGpng grab.E12777.png r1 manage 23.8 K 2009-03-23 - 07:23 SandrinaFernandes Effect of increasing from 3 to 5muA in the resistivity_target
PNGpng grab.y11880.png r1 manage 22.7 K 2009-03-23 - 02:09 SandrinaFernandes Effect on resistivity when increasing proton beam from 1 to 3 muA
PNGpng image001.png r1 manage 105.2 K 2009-03-25 - 17:59 ThierryStora T°300AorW heating Stef
Microsoft Word filedoc picture1.doc r2 r1 manage 30.7 K 2009-03-22 - 08:28 SandrinaFernandes T_vs_Iprotons
Microsoft Word filedoc report-672335.doc r1 manage 43325.0 K 2010-07-12 - 21:19 UnknownUser  
Edit | Attach | Watch | Print version | History: r51 < r50 < r49 < r48 < r47 | Backlinks | Raw View | WYSIWYG | More topic actions
Topic revision: r51 - 2020-08-18 - TWikiAdminUser
 
    • Cern Search Icon Cern Search
    • TWiki Search Icon TWiki Search
    • Google Search Icon Google Search

    Sandbox/SandboxArchive All webs login

This site is powered by the TWiki collaboration platform Powered by PerlCopyright &© 2008-2023 by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
or Ideas, requests, problems regarding TWiki? use Discourse or Send feedback